### v3

# Generator-Based Fuzzers with Type-Based Targeted Mutation

Abstract
Generator-Based Fuzzers (GBF) is a promising technique that allows fuzzers to construct inputs using programmable generators. This approach can be particularly useful for fuzz applications with complex structured input as it encodes the syntactic structure of the application’s input within its input generators.
However, for a large complex input using a uniform mutation strategy to construct new inputs can have limited effectiveness. This is because, the most valuable locations to mutate within the input shrink as a fraction of its size. This suggests that, as with any fuzzer, directing GBF to reach particular code targets can increase the its effectiveness. In previous work, coverage-guided fuzzers used a mix of static analysis, taint analysis, and/or constraint-solving approaches to address this problem. However, these techniques are unlikely to scale to modern applications. This is because tracking the complete data flow in a cross-language and distributed environment as found in modern applications, e.g., serverless computing, would be very difficult.
In this paper, we introduce a type-based mutation heuristic for Java GBF. Our key intuition is that if one can identify which sub-part (types) of the input will likely influence the branching decision, then focusing on mutating the choices of the generators constructing these types is likely to achieve the desired coverage. We used our technique, along with constant string lookup, to fuzz AWS Lambda applications. Results compared to a baseline GBF tool show an almost 20% average improvement in application coverage, and larger improvements when third-party code is included.
Index Terms:
Testing and Debugging, Program analysis, Testing tools.
I Introduction
Fuzzing (i.e., large-scale automated random testing) can be one of the most practical ways to find bugs that might otherwise be missed by developers, However, even given recent progress, fuzzing is limited in the scale and complexity of modules to which it can effectively be applied. Two techniques, coverage feedback and input generators, help make fuzzing more practical, but we observe a missed opportunity when they are used together for larger inputs and codebases.
Coverage-guided fuzzing (hereafter “CGF”), as popularized in AFL [1], augments mutational fuzzing by recording which code locations a test input covers. CGF then prioritizes inputs that cover previously-uncovered code (intuitively, “interesting inputs”) as the starting points for future mutations. CGF is effective at biasing the search process towards relevant inputs without requiring expensive analysis, and this power has made it popular both for practical bug-finding and as the basis for research on further enhancements [2, 3, 4, 5, 6, 7, 8]. However CGF does not imply any change to the mutation process, which is still random.
Another idea that can improve the effectiveness of fuzzing is to capture patterns of relevant input in code with input generators. An input generator for a given type (also commonly used in property-based testing such as QuickCheck [9]) is just a function that returns a random instance of the type. Using generators to provide inputs yields generator-based fuzzing (hereafter “GBF”). GBF can be used for both unit and system testing, and can capture patterns that constitute legal and typical inputs. The effort of writing generators can be amortized because many data types (e.g., associated with common file formats or libraries) are reused across applications, and generators for composite types can be expressed in terms of generators for their components. More recently, the authors of the Zest system [10] demonstrated how CGF and GBF approaches can be combined by applying mutations to the sequence of random choices a generator uses to build an object. Such mutations are more likely to achieve semantically relevant coverage because the generators enforce input validity constraints. However even when CGF and GBF are combined as in Zest, the location for a mutation is still chosen uniformly at random across the input. We observe a missed opportunity that such tools still have no connection between their coverage goals and the process of selecting mutations.
More precise targeting of mutations becomes increasingly important as test inputs become larger. On a small input, choosing mutation locations uniformly at random does not require much time to cover every part of an input, just by brute force. But for larger inputs, the most valuable locations to mutate shrink as a fraction of the input size, such as if a format’s metadata has a fixed size while the data it describes gets larger. For large inputs, selecting mutation locations uniformly at random can be wasteful.
Similarly, being selective about relevant code becomes more important when fuzzing larger programs that make extensive use of libraries and frameworks. Libraries and frameworks improve productivity what they encapsulate rich functionality and allow the code that uses them to be relatively simple and high-level. But in this case, the code newly written for an application may be only a small fraction of the total code executed. We will use the term “application” to refer generically to the often smaller code that was written for a specific use case and that is the primary target of fuzzing. We will use “third-party code” to refer generically to all of the other code that also executes as part of the program under test but is not the primary target of fuzzing, such as libraries and frameworks. If a fuzzing tool does not distinguish between coverage in the application code and third-party libraries, it may spend most of its time exploring the third-party code. This can lead to an illusion of progress, as increased coverage in the third-party library does not necessarily translate to meaningful coverage in the application code.
For instance, in a program that performs image processing, like the one we present in more detail in Section II, a small amount of code in the application makes decisions based on the image metadata such as its size and creation time. But the metadata is small compared to the image’s pixel data, and the main program is small compared to the third-party image I/O and image processing libraries. To scale fuzzing as file size and third-party code usage become significant, fuzzers must concentrate their effort to exercising branches in the application by targeting the most relevant parts of the input.
For our main evaluation, we fuzz software for serverless microservices written in Java for AWS Lambda and compatible platforms. This is a practically important area where it is common for a small piece of application software to depend on a complex stack of third-party libraries and frameworks. The serverless domain also illustrates why it is appealing to use a lightweight fuzzing technique that does not require entire data flows and software stacks to be instrumented. More expensive analysis techniques such as taint tracking or symbolic execution could also be used to trace which parts of the input are relevant to code [11, 12, 13], but tracking the complete data flow in a cross-language and distributed environment as found in serverless computing would be very difficult.
The core idea of our approach is to link input generation to code in the application that uses that input via types. Part of this link happens statically and another happens dynamically. In the static phase, we use a simple analysis that determine for each branch in the program under test a ranked list of influencing types of objects that are data-flow predecessors of the branch condition. This information is later used to guide the fuzzer toward uncovered application branches, by prioritizing mutating choices within input generators responsible for constructing objects of these statically identified influencing types.
In the dynamic phase, we associate types to dynamically constructed objects from the input generators. Since generator methods commonly call other generator methods to construct complex types, our implementation uses a simple form of execution indexing [14] to record the type information for input-generating code in the context of its call stack.
Finally, to achieve effective mutation, our technique matches types statically associated with uncovered branches, to types dynamically associated with parts of the input. Our technique increases the frequency of mutating parts of the input corresponding to objects of matched influencing types. As in other forms of CGF, this mutation occurs in a loop that measures whether mutated inputs achieve new code coverage, and saves inputs that trigger new coverage to be a basis for future mutations. As an additional improvement to accelerate the fuzzing process, our static analysis also collects constant strings from the application code, and includes these strings as possibilities when generating strings in program inputs.
We evaluated our technique on AWS Lambda applications, as they demonstrate the challenge of fuzzing application code that has a large third-party codebase and loosely defined, but complex and structured, inputs. The results show that our technique improves the code coverage achieved by fuzzing. Coverage of application branches improved by an average of 18.2% (geometric mean) over the baseline, and coverage including third-party code increased by a geometric mean of 43.2% when the set of covered libraries remained the same, and even more when our approach covers additional libraries.
The paper is organized as follows: Section II provides an example, Section III outlines our technique, Section IV presents our evaluation, Section VI discusses related work, and Section VIII concludes.
II Motivating Example
(a) Program for creating thumbnails for images modified in 2008.
(b) Helper method for computing the getting modification year.
Figure 1: A program snippet that creates thumbnails for images of a specific year.
Consider the program in Fig. 1(a). The program creates a single image composed of thumbnails of all images modified in the year 2008. Given the location of image files in inputPath (line 7), the program first finds all files within the specified path (line 9), then it iterates over all image files within this directory to collect those whose modification date matches the intended year (lines 11-13). The helper function yearTaken (in Fig. 1(b)) finds the system modification year of a file object. The method first attempts to read attributes of the file system for the image path, particularly the last modified time (line 2-4). Then, it converts that time into a date form captured in a GregorianCalendar object (line 5-6). Finally, the method extracts and returns the year of the last modification of the file (line 7). The thumbnail program in Fig. 1(a) creates a combined thumbnail image for the filtered images (line 15), using some image processing happening within the regroupInSingleImage method, whose details are omitted.
Assuming that the fuzzer knows that the input to the program includes all the images in the identified folder, the fuzzer would still have difficulty covering line 13. This is because while the meta-data part of the image files is much smaller than the content of the images and obviously important for achieving coverage of line 13, the fuzzer’s random mutation strategy does not favor modifying the metadata portion of the input.
II-A Generator-Based Fuzzers Background
Figure 2: Generator-Based Fuzzer - FCI is the fuzzer chosen input
To fuzz a program, GBF’s fuzzing driver works as follows (Fig.2). First, the fuzzer driver starts by creating program inputs. This step happens indirectly by running the input generators on the Fuzzer-Chosen Input (FCI), see Fig.2. The FCI captures the random values used by the input generators to construct a random program input. The FCI is consulted whenever the input generator makes a pseudo-random number call. This design allows the GBF to replay previous program inputs, by reusing the same FCI sequences, or create new program inputs by mutating subsequences within the FCI; i.e., using the mutation process. Next, the program under test is executed on the constructed program input, while collecting coverage and output information. Useful inputs, i.e., inputs that achieved new coverages, are saved act as the base of mutation in subsequent fuzzing.
Program inputs generated by this approach are more likely to explore deeper parts of the program under test than those created by traditional CGF. This is because generated program inputs are structurally correct, enabling them to bypass the program’s well-formedness syntax checks. However, because it relies on a uniform random mutation strategy to generate new inputs, finding the right combination to trigger interesting program behavior becomes increasingly difficult as the input space increases.
(a) Snippet of the file input generator. Invocations of enclosed generators are highlighted in grey.
(b) Snippet of the calendar generator. Random parametric choices controlled by GBF are highlighted in grey.
Figure 3: A snippet of generators
II-B Fuzzing using GBF
Zest [10] is the state-of-the-art GBF for Java programs. Fig. 3 shows a snippet of the generator’s code written with Zest that constructs a file using a file generator FileWithAttributesGenerator that uses, among other generators, the calendar date generator CalendarGenerator Fig. 3 to generate dates. Both classes define a generate method that constructs the object of the intended type. For example, the FileWithAttributeGenerator’s generate method starts by creating a random filename using the StrGenerator (line 4), then it creates an instance of the FileGenerator that generates the file’s contents (lines 5,6). The remaining code generates random dates (using the CalendarGenerator) to set the creation, the last access, and the last modification dates (lines 9-14). The To create a calendar object, the GregorianCalendar generator choose three random integers for its day, month and year. The year is generated using a geometric distribution with a mean of 5, making it more likely to construct a year that is within $\pm$5 years range from the current year (25%-75% future and past years, respectively) (lines 26-29).
Observe that Zest does not directly control the program input. Instead, it controls, i.e., providing an implementation for, the random selection within the input-generators represented with operations on r, the SourceOfRandomness, highlighted in grey. All such selections is what we refer to as the Fuzzer Controlled Input (FCI). By controlling these random selections, the GBF can indirectly control the generation of program inputs. For example, the GBF fuzzer will control the random selection happening on line 5-7 in Fig. 3.
If one would classify the content of the FCI based on the types they are used in generating, we can see a classification that is similar to the one shown in Fig. 4. The entire FCI is used to generate an object of a File type, which can be further decomposed to subsequences that were used to generate an object of a BufferedImage type (the payload of the image), and the other part is used to generate the meta-data of the file (String and GregorianCalendar types to generate the name and the date information of the file).
The image payload occupies the majority of FIC because it is significantly larger than the metadata. As a result, the fuzzer must make far more random selections to construct the payload than the metadata. For instance, if GBF generates a 100x100 image where each pixel requires red, blue, and yellow color values, the payload segment in FCI would contain 100x100x3 data points. In contrast, the metadata segment would only contain 26-30 data points. This imbalance in the composition of the FCI introduces a bias in mutation, that is, the random uniform mutation strategy will likely introduce modification to image’s payload than its metadata. Consequently, mutations often fall outside the relevant subsequence corresponding to the file system attributes needed to achieve coverage in line 13 in Fig. II. In fact, in a small experiment of an hour long, with 20 repetitions, results show that the baseline FCI was only able to cover line 13, 35% of the time (7 out of 20), while our fuzzing extension achieved the same coverage with a success rate of 70% (14 out of 20).
II-C Fuzzing Using The Proposed Type-Based Mutation
Making targeted mutations necessitates understanding which parts of the FCI need modification. Directing GBF fuzzers is especially challenging because of the lack of mapping between the FCI and the program’s input. Although heavyweight techniques such as taint analysis, symbolic execution, or certain dataflow analyses can be employed to trace the propagation of the FCI within the program, this process is often complex and resource-intensive, potentially impacting the fuzzer’s performance. Additionally, there are scenarios where conducting such analyses may not be feasible due to the presence of heterogeneous languages and systems. For example, it is not easy to track FCI that is going out to a file system.
In this work, we develop a new type-based mutation heuristic that directs mutations toward likely useful subsequences within the FCI, based on the types of objects they are used to generate. Our technique consists of three main steps. First, it statically identifies types that influence branch decisions in the application code. Then, while generating program inputs, it dynamically annotates subsequences within the FCI with the types they contribute to. Finally, it increases the chances of introducing mutations to subsequences whose types influence the decisions of uncovered branches.
To collect influencing types, our technique starts by identifying code targets for the fuzzing process. A code target is defined as the then-side or the else-side of a branching condition within the application code. For example, the then-side and the else-side of the if-statement at line 12 in Fig. 1(a) are two code targets. To cover the then-side of line 12, we statically compute types flowing into a branch. We call this step influencing type analysis which employs a variation of a def-use data flow analysis. This analysis will then associate {GegorianCalendar, File} as the list of influencing types that affects the decision of that code target.
Next, during the construction of the program input, our technique map every value within the FCI to the dynamic type of objects they are used in generating.
Figure 4: FCI used in generating a File from FileWithAttributesGenerator with our extension of Type annotation
For example, Fig. 4 shows the FCI used in generating a File from FileWithAttributesGenerator, annotated by the types they are used in generating. Keeping such a type structure helps us direct the mutation to types that are likely to affect decisions of uncovered code targets. For example, suppose $v_{1}=16$ and $v_{2}=8$ are parts of FCI used (lines 23, 24) in Fig. 3. Then, one can identify that the dynamic types that $v_{1}$ and $v_{2}$ are part of the GregorianCalendar and File types. These type-annotations can be computed by looking at the generator’s callstack at the point of generating a random number for these expressions.
Finally, during mutation, our technique increases the likehood of changing subsequences within the FCI that corresponds to influencing types of uncovered branches. The closer an influencing type is to a code target the more likely it is choose for mutation. For example, our technique would prioritize mutating subsequences used in constructing GregorianCalendar type, since it is the closest influencing type to the code target at line 12.
We ran the baseline fuzzer and our extension over the program in Fig. 1(a). Each run is an hour long and we repeated the experiment 20 times. Our results show that the baseline fuzzing tool was able to cover line 13 35% of the time (7 out of 20), while our fuzzing extension achieved the same coverage with a success rate of 70% (14 out of 20).
III Technique
input : A generator $g$: $I_{c}\rightarrow I_{p}$
[…]
0 Type: Set of random values $V$
0 Type: Set of types $\mathcal{T}$
0 Type: Set of execution indexes $EI$
0 Type: Set of fuzzer-controlled inputs $I_{c}=V^{*}$
0 Type: Set of fuzzer-controlled inputs $I_{c}=(V,EI,\mathcal{T}^{*})^{*}$  # replacing line 4
0 Data: Set of successful inputs $\mathcal{S}\subseteq I_{c}$
0 Data: Set of failed inputs $\mathcal{E}\subseteq I_{c}$
0 Data: Set of total coverages $\mathcal{C}_{total}$
0 Data: Set of all code targets $\mathcal{K}$
0 Data: Set of uncovered code targets $\overline{\mathcal{C}}_{\mathcal{T}}\subseteq\mathcal{K}$
0 Data: Map of type distance $\mathcal{D}$: $\mathcal{T}\rightarrow\mathbb{N}$
0 Data: Map of target type distance $\Gamma:\mathcal{K}\rightarrow\mathcal{D}^{*}$
10 $\Gamma\leftarrow$ analyze($p$)
2 $\overline{\mathcal{C}}_{\mathcal{K}}\leftarrow$keys$(\Gamma)$$\mathcal{D}\leftarrow$ unifyTypes$(\Gamma)$
3 $\mathcal{S}\leftarrow$ RANDOM } $\mathcal{E}\leftarrow\emptyset$$\mathcal{C}_{total}\leftarrow\emptyset$
4 repeat
5      for $i_{c}$ in $\mathcal{S}$ do
6           for $1\leq j\leq$ numCandidates$(i_{c})$ do
7                $i_{c}^{\prime}\leftarrow$ mutate $(i_{c})$
8                $i_{c}^{\prime},\mathcal{D}^{\prime}\leftarrow$ mutate $(i_{c},\overline{\mathcal{C}}_{\mathcal{K}},\mathcal{D})$  # replacing line 19
9                $i_{p}\leftarrow g(i_{c}^{\prime})$
[…]
11                if result=Failure then
12                    $\mathcal{E}\rightarrow\mathcal{E}\cup\{i_{c}\}$
13                     else
14                          if coverage $\nsubseteq\mathcal{C}$  then
15                              $\mathcal{S}\leftarrow\mathcal{S}\cup\{i_{c}\}$
16                               $\mathcal{C}_{total}\leftarrow\mathcal{C}_{total}\cup\text{coverage}$
17                               $\overline{\mathcal{C}}_{\mathcal{K}}\leftarrow\overline{\mathcal{C}}_{\mathcal{K}}\smallsetminus\text{coverage}$
18
19
20
21
22                               until given time budget expires
return $g(\mathcal{S}),g(\mathcal{F})$
Algorithm 1 Generator-based fuzzer algorithm, modelled after Zest [10], and amended with our extension in grey
Figure 5: Overall Process of Type-Based Targeted Mutation
Fig. 5 outlines the general process of our technique. First, we collect the influencing types for all branching conditions in a static process. Then, during fuzz testing (the dynamic phase), we annotate segments within the FCI with the types of objects they contribute to generating. Then, the targeted mutation process uses the influencing types information to mutate segments of the FCI with a matching dynamic type. In the process, coverage is collected, and influencing types are updated to reflect their selection priority. Additionally, we enable using of existing string constants, collected during the static phase, as string values for string generators, as they usually present some magic values within the program under test. In the rest of this section, we present the formal algorithm for the GBF extended with our changes for the type-targeting purposes, then we discuss in detail each process within the static and the dynamic phases.
III-A Algorithm
Algorithm  1 shows the main algorithm of GBF with our extension highlighted in grey. We start by describing how GBF works, then we will introduce our extensions. In general, GBF’s algorithm takes two inputs: an input generator $g$, and a program $p$. A generator is a function that given a FCI $I_{c}$, it outputs a program input $I_{p}$. Program $p$ is tested on an input from $I_{p}$, then terminates with either a Failure or a Success result [^1] We assume that $p$ is always terminating. In the implementation a timeout is imposed for each test.. ${V}$ represents the set of random values, the set of fuzzer-chosen input $I_{c}$, which is defined as a sequence of random values. $\mathcal{S},\mathcal{E}$ represent the set of successful and failed inputs, respectively. And finally, $\mathcal{C}_{total}$ represents the set of all code coverage.
GBF proceeds as follows. First, the set of successful inputs is initialized with a single randomly chosen FCI (line 15), while the failed inputs and the total coverage sets are initialized to the empty sets (line 15). Then, the fuzzing loop proceeds until a specified time budget expires (lines 16-30). At each step, an input $i_{c}$ is selected from $\mathcal{S}$, and the number of candidates to generate from $i_{c}$ is determined using the numCandidates heuristic (line 18). This function decides how many mutants will be produced from $i_{c}$. While the specifics of numCandidates are omitted, the number of mutants generally depends on how interesting $i_{c}$ is–specifically, how many new coverages are seen when executed. In each iteration, the selected input is mutated to generate $i_{c}^{\prime}$ (line 19). Mutation is done by introducing random changes to random locations within $i_{c}$. Next, the generator is then invoked on the mutated inputs (line 21) to create program inputs $i_{p}$, which the program executes (line 22). The fuzzer collects the result of the test and the coverage information (line 22). If the test fails, the program input $i_{c}$ is added to the set of failed inputs (lines 23-24). Otherwise, if new coverage was obtained, then GBF saves $i_{c}$ in the successful inputs and total coverage is updated (lines 26-29).
Our extension introduces $\mathcal{T}$, which represents the set of all possible types, and the set of execution indexes $EI$, which captures the unique dynamic location of a particular execution point. Thus, the fuzzer-chosen input $I_{c}$ (line 5) is defined as a tuple consisting of the randomly generated values $V$, the execution indexes $EI$ that specify the unique dynamic execution location within the input generators at the moment a specific random value is generated, and the sequence of types $T^{*}$, which records the return value of each method in the call stack at the point of generating the random value.
In addition to that, our extension introduces four additional data structures: (1) A set of all code targets ($\mathcal{K}$), (2) the set of uncovered code targets ($\overline{\mathcal{C}}_{\mathcal{K}}$), (3) the map of types to their distances ($\mathcal{D}$), and (4) the map of code targets to their typing distances ($\Gamma$). Note that, the distance of a type indicates how far a type is from flowing into the code target. The intuition is that, the closer the distance of a type the more likely it is affecting the branching decision, and thus it is more likely to be selected by the fuzzer for mutation.
Before fuzzing starts, we analyze the program $p$ to collect influencing types together with their distances in $\Gamma$ (line 13), which initializes the set of uncovered code targets. Next, since a type can appear multiple times during the analysis with varying distances to different code targets, we unify their typing typing distances (line 14) by associating a single distance to each type.
Our most significant is happening on line 20. Here, our extension uses a type-based mutation strategy. This process takes the updated fuzzer-chosen input $i_{c}$, the uncovered code targets $\overline{\mathcal{C}}_{\mathcal{K}}$ and the map of typing distances $\mathcal{D}$. The result of this operation is a mutated input ($i_{c}^{\prime}$), and an updated typing distance map $\mathcal{D^{\prime}}$ (line 20), which can de/prioritize types depending on whether a statically identified types was matched with a dynamic one. Finally, GBF proceeds by running the program on the newly constructed program input. If new coverage is achieved then the newly covered code target(s) is removed from the set uncovered code target (line 29). We discuss each component in detail in the following sections (III-B - III-C).
III-B Static Phase
1.v17 = arrayload v10[v26]
2.v19 = invokevirtual File.toPath()Path v17
3.v21 = invokestatic YearThumbnails.yearTaken(Path)I v19
4.if(ne,v21,v3)
Figure 6: Simplified Wala IR for lines 12 in Fig. 1(a). v10,v17,v3, and v26 represent imageFiles, file, the value 2008 and the length of imageFiles
1.(Ljava/util/GregorianCalendar, 2)
2.(Ljava/nio/file/attribute/FileTime, 4)
3.(Ljava/nio/file/attribute/BasicFileAttributes, 5)
4.(Ljava/nio/file/Path, 6)
5.(Ljava/io/File, 8)
6.([Ljava/io/File, 9)
Figure 7: A simplified snippet of the influencing types found for the branch at line 13 in  Fig. 1(a). Each line shows the computed influencing type and its distance from the decision.
In this step, our extended algorithm analyzes the application code to find all code targets. Using this analysis we can direct the fuzzer to cover the code targets identified in this step. For each code target, we compute types likely to influence their decision. An influencing type is any type that can flow to the operands of the branching condition of the code target.
Our analysis computes a variation of def-use relation among variables. We use the Wala [15] static analysis framework to collect influencing types for this analysis, whose IR is in a static single-assignment form [16] representing the bytecode. For example,  Fig. 6 shows a simplified Wala’s IR representation of line 12 in  Fig. 1(a), where the imageFiles is first loaded. This corresponds to loading elements of the array within the for loop at line 11. At Fig. 6:L2 the image file is retrieved using toPath(), followed by invoking the application method yearTaken() to find the year of the photo (Fig. 6:L3), then finally branching on the resulting output (Fig. 6:L4). Note that our analysis is not field-sensitive, for example, although the image file has three date fields, our analysis only captures the GregorianCalendar as one of the influencing types. The analysis consists of three passes: a variable dependency pass, a type collection pass, and a type unification pass.
Variable dependency pass: In this pass, we use a flow-sensitive analysis where we create a dependency graph between variables using a variation of the traditional def-use dataflow. Our extension gathers variable dependencies through a depth-first traversal of the application’s call graph. To capture the variable dependency when a method invocation is encountered, we add more edges to the dependency graph to capture the dependencies between the actual and formal parameters of the method. Similarly, if there is a return statement within the method, then we create a dependency between the returned variable in the callee and that of the caller. Also, for non-static methods, we capture the def-use dependency between the variables passed to the arguments and the reference variable on which the method invocation is done. Our analysis can identify recursive calls but visits the method only once at each encounter. Note that, for any invocation of a non-application method, i.e., a library call or a Java standard library call, we avoid analyzing their body. For example, the edges (use $\rightarrow$ def) in the dependency graph for this pass for  Fig. 6 are:
$\{{\tt v17}\rightarrow{\tt v10},{\tt v17}\rightarrow{\tt v26},{\tt v19}\rightarrow{\tt v17},{\tt v21}\rightarrow{\tt v19},$
${\tt farg({yearTaken})}\rightarrow{\tt v19},{\tt v21}\rightarrow{\tt ret({yearTaken})}\}$
$\cup{\tt E_{yearTaken}}$
Here, the first line captures the dependencies between def-use variables. The second line captures the dependencies encountered due to the method invocation of yearTaken(), where farg(yearTaken), and ret(yearTaken) are functions that return the variable name of the formal argument as well as the variable name of the return variable of the same method. Observe that v21 also depends on EyearTaken: the set of variable-dependency edges computed from analyzing the yearTaken() method. By contrast, the analysis of the body of toPath() is skipped since it is not identified to be defined within the application code (within the user’s outermost package name).
Also, in this pass, we find and collect the type for each variable encountered. To find types, we use the static type signature existing in the code as well as Wala’s type inference. If a variable was found to have multiple types, we keep all of them. In general, we exclude primitive types and java.lang object types.
Type Collection Pass: In this pass, our extension walks all conditions of code targets and computes their influencing types. This is done by walking the variable dependency graph from the previous step for each operand within the code target branch condition. That is, we traverse the dependency graph backward, collecting encountered variable types and their respective distances. The distance to an influencing type indicates its proximity to the target branch, reflecting its significance to the mutation process. The intuition here is that closer influencing types probably have a larger effect on the branch’s decision. Moreover, the closer the influencing type is, the more likely it is to be fine-grained, which means, that the mutation can now happen over a smaller subsequence within the FCI. For example,  Fig. 7 shows a simplified subset snippet of the influencing types collected for the branch at line 12 in  Fig. 1(a). Each line is a pair of a type and a distance. Note that there might be multiple types at the same distance. In this case, we can see that the influencing type GregorianCalendar is much closer than the File type. This suggests that changing subsequences within the FCI related to the generation of the GregorianCalendar is more likely to make necessary changes to cover both arms of the intended branch. Finally, note that we ignore collecting Java String types, primitive types, and their corresponding wrapper classes, such as Double and Float. This is because these types are too common and do not provide sufficient distinction for targeting specific conditions.
Type Unification Pass: In this pass, we merge all influencing types with their distances for all code targets into a unified structure ($\mathcal{D}$). If influencing types are unique among all code targets then the unified structure will be their union. However, if an influencing type appears in multiple code targets with different distances, then the type is added to $\mathcal{D}$ with its distance in the largest code target; i.e., the code target with the largest number of influencing types. The intuition is that the more influencing types the code target has, the more likely the code target can be covered if its corresponding influencing types are tweaked. For example, if $(t,d_{1})$ and $(t,d_{2})$ are found in two code targets $k_{1}$, and $k_{2}$, then the unified structure will contain ($t,d_{3}$) such that $d_{3}=d_{1},\text{ if }|k_{1}|>|k_{2}|$, otherwise $d_{3}=d_{2}$, where $|k|$ denotes to the number of influencing types associated with a code target. Later, distances of various influencing types are updated during the dynamic fuzzing process depending on whether a static influencing type could be matched with a dynamic type. The change in distances allows the fuzzer to prioritize or de-prioritize influencing types depending on whether they were useful dynamically.
Creating Constant String Lookup Table In this step, we implement a simple analysis to identify magic strings, i.e., string that have specific meanings in the program under test. Our analysis gather all reachable string constants from the call graph’s entry point and within the application code. Later, we enable the fuzzer to alternate between selecting magic strings and generating fresh ones, when generating string values.
III-C Dynamic Fuzzing Phase
In this step, we use information about influencing types collected from the static phase to make type-targeted mutations. Recall that there is no direct connection between the FCI and the program input. To enable type-directed mutation, one must first establish this connection between segments within the FCI and the types of objects they are used in generating. Annotating the FCI with types alone does not guarantee that mutated values will generate the same object type. This is because introduced mutations can alter control flow within generators, meaning that values may not be consumed in the same order, which might result in creating different objects than intended. To ensure consistency, we must associate with each data point in the FCI the dynamic location within the generators where it was used. We do this by using execution indexing. Execution indexing (EI) [17, 18] of a program provides an ordered, unique representation of the dynamic trace points within the execution. Execution indexing technique has many applications such as program alignment [19], and detecting deadlocks and concurrency failures [20, 21]. The main idea of execution indexing is that every program point is associated with a vector of code points that identify the unique dynamic index of a particular execution point. We use EI to ensure that values can only be reused in the same dynamic locations. Next, we outline the process of annotating values within the FCI with types, creating their execution indices, and utilizing this information for type-targeted mutations.
III-C1 Using Execution Index to annotate the FCI with Types
To annotate the FCI with types, we intercept invocations requesting a random number, as in r.nextInt(). Then, we collect the type of objects that are currently being constructed. We do that by collecting the return type of methods within callstack. For example, we intercept the execution of r.nextInt(31) in line 23 in  Fig. 3. Let’s assume that the entry within the FCI for this random selection is the number 9. Then, we associate this value with the return types of all method invocations types in the current callstack. In this case, the value 9 will be annotated with {GregorianCalendar, File} types; i.e., (9,{GregorianCalendar, File}). As previously noted, annotating FCI values with types alone is insufficient for targeted mutation.
In targeted mutation, the goal is to introduce localized changes to objects of a specific type while preserving the rest of the input. To prevent sequential processing of the FCI, which could disrupt this localization, we annotate each value with its unique dynamic location in the code by computing its execution index. For example, consider the case where initially generating a fileName required that the StrGenerator use the first 10 values of the FCI. Suppose a mutation occurred that resulted in having StrGenerator using only the first 5 values of the FCI. In that case, the remaining 5 values from the original FCI, while being annotated with String types, will be used within the FileGenerator to create an object of File type. This is a mismatch between the object types created by the same FCI’s subsequence from one test to another.
To avoid problem we associate an execution index with every value of the FCI. The main idea of the execution index is that every program point is associated with a vector of code points that identify the dynamic EI of a particular execution point. In our extension, we use a relaxed definition of execution indexing [14] implemented in Zest, where the uniqueness property among different execution indices is not guaranteed. More precisely, let $m_{1},m_{2},..,m_{n}$ be the callstack when the execution of an expression $p$ is about to happen. The execution index of $ei(p)$ is defined as a vector of pairs $[l_{i},q_{i}]$ such that ($0<i\leq n$), where $l_{i}$ is the label of the statement within method $m_{i}$, and where $q_{i}$ is the number of times the statement was invoked during the current invocation of $m_{i}$.
To illustrate, consider the input generator at  Fig. 3 where a file is created and its creation date is constructed by invoking the CalGenerator at line 13 ( Fig. 3). If lCal is created during the first iteration of the while loop on line 9, then the execution index of the random choice of selecting a day of a month (line 23 in  Fig. 3); i.e., $ei(23)=[13,1,23,1]$. However, if lCal is created during the second iteration of the while loop then the execution index of the random choice of the day of a month would be $ei(23)=[13,2,23,1]$. This is because the generate(r,..) method at line 13 in this case has been invoked twice in the context of FileWithAttributesGenerator.
Using this approach the FCI ($i_{c}$) is represented as a sequence of triples of the form $i_{c}\in(V,EI,\mathcal{T}^{*})^{*}$. For example, continuing our above example of creating the lCal object (line 13). Our extended FCI for selecting the day of the month (line 23) will be the triple of the form: $(9,ei(23),\{GregorianCalendar,File\})$ Note that this technique is most effective when there is a separate generator method for every level of object that the generator constructs, a common coding practice.
III-C2 Type-Based Mutation
In this step, we use the dynamically captured type distances maintained in $\mathcal{K}$ to select a type to mutate. The type selection process works as follows. First, we sort all types based on their distances, then we normalize their distances. Next, we randomly choose a number and return the first type whose distance is greater than or equal to the selected random number. Then, we match the selected type to those in the FCI. If an EI object is found with the type from the previous step, our algorithm proceeds in two steps. First, it decreases the type’s distance (multiplying it by 3/4). Second, it randomly mutates a subsequence of a random length starting from the matched EI object. If no EI object was found with a matching type, then the type’s distance is updated (multiplying it by 4/3) to decrease the type’s probability of being selected in future iterations.
IV Evaluation
Our research questions are:
TABLE I: List of AWS lambda applications obtained from GitHub with the number of enclosed lines of code (loc)
Serial benchmark loc
1 s3-java 174
2 anishsana 101
3 load-historic-data 182
4 lambda-unzip 170
5 csv-loader 407
6 nikoshen 131
7 upload-survey 16,661
RQ1:
Does the SpotOn achieve more coverage than GBF?
RQ2:
What is the effectiveness of the type-based targeting without the constant string optimization?
RQ3:
What is the overhead of the type-based targeted mutation?
We evaluated our technique using AWS Lambda applications. These applications tend to be compact in the size of user-written code while using extensive code from third-party libraries. Additionally, their inputs often possess a degree of complexity and structure, although loosely defined. We collected our benchmark suite from GitHub. We searched for AWS Java Lambdas that are triggered by an S3EventNotification and that use a single S3, and/or a single DynamoDB service. We removed applications with no branches and those that use unsupported formats.
Tab. I lists of benchmarks we used [22], which include:
  1. 1.
    s3-java [23]: when an image file is inserted/updated, the lambda scales it and puts it back in the bucket.
  2. 2.
    anishsana [24]: when a CSV file of students’ grades is inserted/updated into an S3 bucket, the lambda computes an average grade and places a record into a DynamoDB table.
  3. 3.
    load-historic-data [25]: when a stock ticker CSV file, the lambda collects some data from the file and creates a corresponding record onto a DynamoDB table.
  4. 4.
    lambda-unzip [26]: an archive is inserted/updated into an S3 bucket, the lambda unzips its content into the bucket and removes the archived file.
  5. 5.
    csv-loader [27]: when an archived CSV file is inserted/updated into an S3 bucket, the lambda it extracts the CSV file; and creates a DynamoDB entry for each record within the CSV file.
  6. 6.
    nikoshen [28]: when an object is inserted/updated into an S3 bucket, the lambda it checks whether a matching DynmoDB item exists. If not, it creates it.
  7. 7.
    upload-survey [29]: when an object is inserted/updated into an S3 bucket, the lambda serializes the received data, and updates the DynamoDB with their information.
-Implementation: We used Zest [10, 30], the state-of-the-art Java fuzzing tool to add our extension. Zest is a GBF that allows users to write input generators. To compute coverage, Zest instruments dynamically loaded classes to collect coverage information. We built our extension on top of Zest; we call our tool SpotOn [22]. Our implementation utilizes an experimental codebase within Zest that defines the FCI as a set of execution indices with values. We extended this work to amend the types of the generator’s call stack to each execution index object of the FCI. In both tools, we added a heuristic that controls the number of mutants created from each FCI. The heuristic ensures inputs with significant time are not overrun with excessive children. As processing time for children increases, the number of tests decreases. This heuristic embodies numCandidates$(i_{c})$ in Alg. 1. Additionally, we utilized the Wala [15] framework for static analysis, running it offline before executing SpotOn. Lastly, in SpotOn, mutation alternates between targeted and random mutations with a 50% probability.
TABLE II: Average application coverage, (+SpotOn): only found only by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
benchmark common +SpotOn +Zest
lambda-unzip 21.2 1.80 0
csv-loader 13.45 4.45 0.7
anishsana 8.45 0.50 0.05
load-historic-data 13.15 0.80 0.05
nikoshen 12.60 2.9 0
s3-java 16.65 0.35 0
upload-survey 10.00 6.10 0
-Generators: To fuzz test AWSLambda applications, we wrote an AWS generator package that populates AWS resources interacting with the AWSLambda application, such as S3 and DynamoDB instances. We assume that the structure of the DynamoDB table is known i.e., its primary key and range key; if one exists. Before every test, the entire AWS state is cleared and repopulated. We randomly generate files of various types and add them to an S3 bucket. Supported types include XML, image, CSV, text, Java class, JavaScript, and archive files. Some of these formats already existed in Zest, others were created by the authors of this paper. These generators have a 0.003 probability of creating erroneous files, like missing or incorrect file extensions. In contrast, for DynamoDB, we randomly generate items with various column sizes (excluding primary and range keys) and populate the resource with them.
-Setup: We used LocalStack 1.4.0 [31]: a local cloud software development framework used to develop, test and run AWS applications locally. We ran our experiment over a Dell Inc. Precision T3600 machine with 32 GB RAM running Ubuntu 22.04.4 LTS. In our evaluation, we utilized Zest’s Maven plugin, JQF, which instruments coverage code within dynamically loaded classes for fuzzing. We collected coverage information for condition coverage and method invocation coverage. Coverage of already loaded classes (e.g., java.lang, etc.) is not included in the results of this section. Additionally, coverage for infrastructure applications like com.amazonaws and com.ibm.wala was excluded, as it is beyond the scope of this research.
TABLE III: Average of all (third-pary-included) coverage, (+SpotOn): only found by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
benchmark common +SpotOn +Zest
lambda-unzip 126.8 14.15 1.05
csv-loader 86.3 42.9 3.45
anishsana 15.45 0.5 0.05
load-historic-data 119.5 7.05 1.6
nikoshen 218.1 702.55 0
s3-java 485.5 0.50 0
upload-survey 88 2350.55 0
IV-A RQ1: Does SpotOn achieve more coverage?
To answer this question, we fuzz tested each benchmark using both Zest and SpotOnfor an hour with 20 repetitions. We categorize coverage results into two categories. (1) Application code coverage: defined as coverages of conditional branches and method invocations within the application code. (2) All code coverage: defined as coverages of conditional branches and method invocations within the application code as well as for those happening within all non-excluded libraries.
Tbl. II shows results for the application coverage, where Common lists the common code targets found by both tools, while additional ones found by SpotOn, and Zest are listed in +SpotOn, and +Zest, respectively. The (*) indicates the statistical significance of the coverage, which shows that the results of 4 out of the 7 benchmarks’ results are statistically significant refuting the null hypothesis that the difference in the coverage results between the two tools is only due to random variation.
Tbl. III shows results for all-code coverage, showing that the coverage of three out of the seven benchmarks is statistically significant. In more detail, we notice that in both tables, most of the code targets were found by both tools, though SpotOn on average finds more code targets. This is mostly apparent in nikoshen and upload-survey. In nikoshen type-targeting was useful to focus the mutation on the DynamoDB Item to create a matching and a valid item. In upload-survey type-targeting was useful to focus on the name of the S3 object. Both benchmarks used the constant string lookup to generate specific string constants. In other benchmarks, we notice that type-targeting is useful for more stable results (more in the following section).
Finally, observe the correlation between application coverage and overall code coverage. Achieving higher application coverage generally leads to increased overall coverage. This is because covering a new branch in the application code often triggers the execution of additional code from third-party libraries. Thus, prioritizing application coverage is likely more efficient and leads to higher overall code coverage
IV-B RQ2: What is the effectiveness of type-based targeting?
To isolate the effect of the type-based targeted mutation from the constant string optimization, we separated the later into a separate mode: ZestStrOpt. Tbl. 15 through Tbl. 15 shows the overall coverage for Zest, ZestStrOpt, and SpotOn. The solid line in the graph describes the median for all 20 repetitions of the experiment. While the shaded lines describe the 95% confidence interval.
The first observation is that in all benchmarks, SpotOn performed either the same or better than Zest. In the case of s3-java, neither ZestStrOpt nor SpotOn demonstrate significant coverage improvements over the baseline Zest. However, we believe this outcome inaccurately represents the effectiveness of type-based targeting, primarily due to a limitation in collecting coverage from dynamically loaded libraries. Particularly in this benchmark, the new coverage achieved by type-based mutation stemmed from pre-loaded classes, which were consequently omitted from the overall coverage collection. This limitation can be circumvented by instrumenting the code prior to the loading of any of these classes, specifically through the use of a Java Agent. Implementing this approach is a task for future work. In lambda-unzip (Tbl. 15), load-historic-data (Tbl. 15), and anishsana (Tbl. 15), while we do not see a significant coverage difference, we can observe that SpotOn results are more reliable among the three tools. This is apparent due to its relatively smaller confidence interval. Similarly, we can see stable results for SpotOn in nikoshen, csv-loader and csv-loader. We see significantly better performance for SpotOn and ZestStrOpt in nikoshen (Tbl. 15), and in csv-loader (Tbl. 15). In fact, the median of the baseline Zest is at the very bottom suggesting that most of the runs were performing poorly compared to SpotOn and ZestStrOpt. The orange shading in csv-loader suggests that a run or two from the population, Zest was able to craft the right input that achieved the desired coverage. In nikoshen the string constant was needed to create a matching state between the DynamoDB item and the S3 object. In csv-loader the string constant was needed to create an S3 object with a particular name. In both benchmarks, the type-based targeted mutation was useful to speed up the coverage by focusing the mutation on the substructure with the matching influencing type. In general, we can conclude that SpotOn has the most reliable performance in coverage finding among the three extensions.
Figure 8: Median coverage of s3-java
Figure 9: Median coverage of lambda-unzip
Figure 10: Median coverage of load-historic-data
Figure 11: Median coverage of anishsana
Figure 12: Median coverage of nikoshen
Figure 13: Median coverage csv-loader
Figure 14: Coverage of upload-survey
Figure 15: Avg. gen. per test; errorbars between $25^{th}-75^{th}$ percentiles
IV-C RQ3: What is the overhead of SpotOn?
To identify the performance bottleneck for SpotOn, we divided the fuzzing operations into four processes:
  1. 1.
    mutation: the process of mutating subsequences of FCI,
  2. 2.
    generation: the process of constructing an input by running the input generators on the FCI,
  3. 3.
    testing: the process of running the program under test with the input generated from the generation step,
  4. 4.
    handling: the process of handling test result, i.e., new coverage, success or failure.
Our primary experiment’s data revealed that the combined impact of mutation and handling processes accounts for no more than 3% of the total execution time. To assess the effect of the generation and testing steps, we conducted a controlled experiment in which SpotOn managed two distinct FCI types: a linear FCI to mimic Zest’s operations and an EI-enabled type-annotated FCI. We utilized the linear FCI for the generation phase and the annotated FCI for the mutation process. This setup ensures that both Zest and SpotOn produce the same inputs, but SpotOn additionally creates and maintains the annotated FCI. We ran both tools using this configuration for a fixed number of tests. For each benchmark, we used three-quarters of the tests from the one-hour time-out experiment. The following results report the averages of the 10 runs.
Tbl. 15 shows the average generation time per test. We notice that the average generation cost of SpotOn is not too far off from Zest’s. This overhead is due to computing and maintaining the annotated FCI. More precisely, during the generation SpotOn intercept, every method invocation to keep track of EI counters used in their creation. Then, when a java.util.Random method is about to execute, SpotOn creates the EI instance and looks up its corresponding value within the FCI. If found, the generation uses the present value, otherwise a fresh value is created and the triple of the EI object, its annotated type, the fresh value is added to the FCI.
Also, we measured the usage of the Java heap during experiments. We noticed that both Zest and SpotOn have a high rate of heap allocations-deallocations forming a sawtooth-like shape for heap usage over time. Overall, we observed that SpotOn’s memory overhead ranges from a minimum of 0.82x to a maximum of 4.62x, with an average overhead of 3.29x when compared to Zest’s.
V Threat to Validity
The effectiveness of the GBF depends on its generators, and so our results are dependent on the generators that we have constructed. We created these generators because there are no previously written generators for the types we used in our experiments, i.e., AWS-related types. In general, the results can be dependent on what the generators produce and how they are defined. The former is an inherent limitation in GBF, and we addressed it by designing generators to produce random inputs for all decision-relevant data typically used in programs while keeping others, like eventTime and eventVersion in S3Event, constant.
The latter limitation can arise where the linear reuse of values in the FCI may unintentionally alter the structure of created objects (see [32]), potentially impacting Zest’s results. This limitation is mitigated by designing generators that construct objects in a logical order without relying on any specific program under test.
VI Related Work
Coverage-guided fuzzing (CGF) [1, 33, 2, 3, 4, 5, 6] is a random testing technique with lightweight instrumentation for coverage computation. Using the coverage information, fuzzing tools can then make informed decisions about the choice of interesting input likely to achieve new coverage if mutated. CGF revealed many bugs in widely used/tested programs, such as in Clang, OpenSSH, JavaScriptCore, LibreOffice, Python, SQLite, Google closure-compiler, JDK, Mozilla, and BCEL to mention a few [1, 34, 35].
However, as CGF lacks information about the input structure, its effectiveness can be negatively affected. Grammar-based fuzzers such as CSmith [36], jsfunfuzz [37], Grammarinator [38], and others [39, 40, 41, 42, 43] use a declarative form of the input that defines input’s grammar. Some works used the existence of a test corpus to craft new inputs from existing tests [43, 44, 45, 46]. Most of this work relied on understanding grammers and/or features of existing corpus and then use structural mutation to generate new inputs.
Generator-based fuzzers (GBF) utilize programmatically defined generators to create structured inputs. While both grammar-based fuzzers [39, 47, 42] and generator based fuzzers [10, 48, 49, 50] are more likely to trigger semantic behavior within the program under test, due to their knowledge of the structure of the input, we view them as alternative techniques to create a structure, semantically correct input. However, we observe that the latter naturally imposing language-type structure, which we leverage to focus the generation on likely useful substructures within the input. Padhye el. [10] introduced generator-based fuzzers by extending generator-based testing [9] to use random input generators. In our work, we extend Padhye el.’s work to enable type-based targeted mutation.
In CGF, obtaining useful seeds has been an active area of research [2, 3, 51, 52], where some work focused on developing heuristics for picking seeds, and others focused on finding values to create good seeds. For example, Marcel Bohme et al.[2] used a Markov chain to assign energy to seeds, prioritizing inputs that exercise low-frequency paths for better coverage. AFLGo[3] computes the closeness of a seed to targeted code locations and gradually assigns more energy to seeds closer to the target. While these approaches focus on input selection heuristics, our work concentrates on mutating inputs based on their types. $\mu^{2}$[51] controlled seed selection by computing a mutation score for each seed, while Pythia[53] used a learning-based technique to identify useful mutations. FairFuzz [54] computes rare test targets and freezes parts of seeds that contribute to reaching these targets, mutating the remaining parts. Other tools[55, 12, 13] use a form of symbolic execution to compute useful values for the input. The key advantage of this technique is that, unlike the heuristic-based approach, symbolic execution can turn the search for new inputs into a constraint set. When these constraints are successfully solved, they yield a new input that effectively covers the intended targets. The cost, however, is that these tools are usually slow due to the overhead of constructing and solving constraints.
The work we present in this paper falls into the first category of heuristics for picking seeds, which tries to craft a likely useful input using type-based targeted mutation. In contrast with FairFuzz, our work finds the places where mutations are needed, rather than the places that should be avoided. More recent work in GBF [49, 32] highlighted the importance of identifying the structure of the input to increase the effectiveness of the mutation. However these techniques focused on orthogonal aspects of mutation, namely, whether change affects the object structure [49], or side-effects of mutations on subsequent choices [32].
VII Future work
One limitation of this approach is its ineffectiveness with recursive types, as the entire FCI will be made up of a few repeated types making the type-based targetting less effective We plan to extend this work to support effective targeting for objects of recursive types.
We focused on testing serverless applications, as they are ideal candidates for our technique. We plan to evaluate our technique on traditional programs. Unfortunately, existing traditional programs rely on using primitive or recursive types, unsuitable for our technique. Thus, we will seek suitable benchmarks for traditional programs to assess our approach.
Finally, we plan to investigate the automatic generation of the generators, a common hurdle for the usability of GBF.
VIII Conclusion
We introduced a novel type-based mutation heuristic with constant string lookup for generator-based fuzzers. Results show improvements in application coverage of an average of 18% over the baseline, and improvements in coverage including third-party code of 43% when the set of covered libraries remained the same, and even more so when our approach covers additional libraries.